OpenStudy (anonymous):

Draw and label a rectangle whose area is represented by the expression. Then write an equivalent in expanded form. 1. (x+1)(x+4) 2. (x+5)(x+6) 3. 3x(5+2)

OpenStudy (danjs):

Area of Rectangle = Length x Width = (x+1) x (x+4)

OpenStudy (danjs):



OpenStudy (danjs):

Thats it, do the same for the other 2

OpenStudy (danjs):

2) Length = (x+5) Width = (x+6) 3) 3x(5+2) Length = 3x Width = (5+2) = 7

OpenStudy (anonymous):

Thank you! :) @DanJS